OpenStudy (anonymous):

Add!: (2a^4-a^3+5a^2+16a-11)+(a^5+8a^3+5a^2-5a+6)+(-5a^4+a^2-6a-4)

OpenStudy (anonymous):

bunch of combining like terms i.e. terms with the same exponent

OpenStudy (callisto):

(2a^4-a^3+5a^2+16a-11)+(a^5+8a^3+5a^2-5a+6)+(-5a^4+a^2-6a-4) =2a^4-a^3+5a^2+16a-11+a^5+8a^3+5a^2-5a+6-5a^4+a^2-6a-4 =a^5+2a^4-5a^4-a^3+8a^3+5a^2+5a^2+a^2+16a-5a-6a+6-4-11 =a^5+(2-5)a^4+(8-1)a^3+(5+5+1)a^2+(16-5-6)a+(6-4-11) =...?